# Neuron Empirical Gradient: Connecting Neurons’ Linear Controllability and Representational Capacity

### v1

@@ Abstract
Although neurons in the feed-forward layers of pre-trained language models (PLMs) can store factual knowledge, most prior analyses remain qualitative, leaving the quantitative relationship among knowledge representation, neuron activations, and model output poorly understood. In this study, by performing neuron-wise interventions using factual probing datasets, we first reveal the linear relationship between neuron activations and output token probabilities. We refer to the gradient of this linear relationship as “neuron empirical gradients,” and propose NeurGrad, an efficient method for their calculation to facilitate quantitative neuron analysis. We next investigate whether neuron empirical gradients in PLMs encode general task knowledge by probing skill neurons. To this end, we introduce MCEval8k, a multi-choice knowledge evaluation benchmark spanning six genres and 22 tasks. Our experiments confirm that neuron empirical gradients effectively capture knowledge, while skill neurons exhibit efficiency, generality, inclusivity, and interdependency. These findings link knowledge to PLM outputs via neuron empirical gradients, shedding light on how PLMs store knowledge. The code and dataset are released [^1] https://anonymous.4open.science/r/NeurGrad.
@@
Connecting Neurons’ Linear Controllability and Representational Capacity
Xin Zhao  Zehui Jiang The University of Tokyo {xzhao,zjiang}@tkl.iis.u-tokyo.ac.jp                        Naoki Yoshinaga Institute of Industrial Science The University of Tokyo ynaga@iis.u-tokyo.ac.jp
@@ 1 Introduction
Although Transformer (Vaswani et al., 2017)-based language models (LMs) benefit from large-scale pre-training, the pre-trained LMs (PLMs) suffer from hallucination, where models generate incorrect knowledge. This issue makes it important to understand the mechanism by which PLMs store knowledge within their parameters (Dai et al., 2022; Niu et al., 2024; Wang et al., 2024a, 2022).
In Transformer-based LMs, feed-forward (FF) layers serve as key-value memory Geva et al. (2021), with neurons possessing the ability to retrieve knowledge. Previous work reveals that specific facts correlate with a limited number of neurons (knowledge neurons) (Dai et al., 2022; Yu and Ananiadou, 2024; Wang et al., 2024b), and even specific neurons own the abilities to perform various language skills (Wang et al., 2022; Tan et al., 2024). Although these studies reveal neurons’ role in handling knowledge and skills, the numerical relationship between neuron activations and model outputs remains poorly understood.
@@
Figure 1: Overview of our contributions: i) observation on the linear controllability of PLM’s outputs by shifting neuron activations, ii) an efficient method, NeurGrad, for computing the gradient of this linear relationship, and iii) skill neuron probing on MCEval8K to confirm empirical gradients capture diverse language skills.
In this study, we first quantitatively analyze how neuron activations affect model outputs through factual knowledge probing (§ 2). To observe model generation under varying neuron activations, we conduct a neuron-wise intervention on PLMs using MyriadLAMA (Zhao et al., 2024), a factual knowledge probing dataset. For given changes in neuron activations, we observe the resulting changes in the probabilities of target tokens for correct knowledge (hereafter, “output probabilities”). Notably, we find that for some neurons, within a certain range of activations, shifts in their activations (hereafter, “activation shifts”) have a linear relationship with the output probabilities. We also find that neurons differ in the direction they shift output probabilities as their activations increase — a property we call polarity, which allows us to classify neurons as either positive or negative. Our evaluation of six PLMs, including Llama2-70B, confirms that neurons generally exhibit both linearity and polarity. We term the gradient of this linear relationship between a specific neuron with a token in response to a prompt as the (neuron) empirical gradient.
While empirical gradients quantify a neuron’s importance and direction in shaping PLM outputs, their calculation is costly due to variability across prompts, neurons, and target tokens. To facilitate quantitative neuron analysis, we thus propose NeurGrad, an efficient method for estimating empirical gradients, and validate its performance on the MyriadLAMA dataset (§ 3). Our results on the above six, diverse PLMs show that NeurGrad outperforms baseline methods in both efficiency and precision.
We then leverage NeurGrad to investigate how empirical gradients represent language knowledge through skill neuron probing (Wang et al., 2022). Different from the factual knowledge probing in § 2, skill neuron probing aims to identify neurons associated with general language skills such as sentiment classification. We create a new multi-choice benchmark (MCEval8K) containing datasets conveying diverse language skills and train classifiers using empirical gradients calculated by NeurGrad as input. Neurons whose gradients provide valuable information for constructing the optimal classifier are identified as skill neurons for specific tasks.
@@
  • We quantitatively confirm that neuron activations in PLMs have linear impacts on output token probabilities, introducing the concept of neuron empirical gradients. (§ 2)
@@
  • We present NeurGrad, an efficient method for estimating neuron empirical gradients. (§ 3)
  • We confirm that empirical gradients serve as indicators of language skill representation via skill neuron probing (§ 4,§ 5); skill neurons demonstrate efficiency, generality, inclusivity, and interdependency (§ 6,§ C).
@@
  • We built MCEval8K, a multi-choice benchmark spanning various skill genres on language understanding. (§ 4.2)
2 Neuron as Linear Output Modulator
In this section, we aim to gain a deeper insight into how neurons in PLMs’ FF layers influence model generations in a quantitative manner. Using factual knowledge probing as the target task, we perform neuron-wise intervention by adjusting neuron activations for the same prompt and observing the resulting change in output tokens’ probabilities.
2.1 Settings
@@
Models. To make the analysis result general, we experiment with two types of LMs, masked and causal LMs, with varied sizes and learning strategies. For masked LMs, we use three BERT (Vaswani et al., 2017; Devlin et al., 2019) models: BERTbase, BERTlarge, and BERTwwm. We construct masked prompts and let the model predict the masked token. For causal LMs, we examine three instruction-tuned LLMs of Llama2 family (Touvron et al., 2023), with sizes of 7B, 13B, and 70B. Following Zhao et al. (2024), we instruct them to generate single-token answers. See § B for model details.
@@
Dataset. We utilize a multi-prompt knowledge probing dataset, MyriadLAMA [^2] https://huggingface.co/datasets/iszhaoxin/MyriadLAMA (Zhao et al., 2024), for neuron intervention. MyriadLAMA offers diverse prompts per fact, reducing the influence of specific linguistic expressions on probing results. We focus on single-token probing, where the target answer is represented by a single token. For each PLM, we randomly sample 1000 prompts from MyriadLAMA, where the model correctly predicts the target token. Due to differences in tokenizers, the probing prompts may vary across PLMs.
@@
Neuron-wise intervention. We conduct neuron-wise intervention to analyze how activation shift affects model outputs. Specifically, we alter the neuron activations within a range of [-10, 10] with a step size of 0.2 to observe the resulting changes in target token output probabilities. Since observing the effect of a single neuron on one token for one prompt requires 100 inference runs and is costly, we only perform the neuron-wise intervention on specific neurons selected by either random sampling and choosing the top-$k$ neurons with the highest absolute computational gradients. [^3] Computational gradient refers to the gradient computed from the computational graph through backpropagation.
2.2 Results and Analysis
From experimental results, we reveal the numerical relationship between neuron activation shifts and output probabilities in PLMs.
@@ 2.2 Results and Analysis
Figure 2: Average absolute Pearson correlation between activation shifts and output probabilities on 1000 neurons $\times$ 10 prompts with a step size of 0.2.
Correlation vs. shift range. We first calculate the Pearson correlation between the shift ranges and the output probability of the correct tokens, considering only the absolute values to examine their linear relationship we call neuron linearity. The correlations are averaged over 10 prompts, each with 1000 neurons, for each activation shift size. [^4] The mean/max/min activations over 1000 prompts on BERTbase are -0.17/4.83/-0.04; On Llama2-7B: -21.6/7.13/0.
Figure 2 depicts averaged correlations for the two neuron selection methods. The top-gradient neurons demonstrate high correlations across the PLMs and shift range, which is higher than the randomly sampled neurons. This suggests that the neuron linearity holds for top-gradient neurons (possibly, knowledge neurons). [^5] The Llama2’s lower correlations for smaller activation shift ranges are due to their small gradient magnitudes. We examine the gradient magnitudes of neurons in the PLMs. As a result, Llama2’s gradients are five orders of magnitude smaller than BERT’s, typically ranging from $10^{-5}$ to $10^{-8}$. Given that gradients are in 16-bit floats with about $5.96\times 10^{-8}$ precision, random noise may overshadow true gradients in correlation calculation on Llama2 with randomly sampled neurons for smaller activation shift ranges. Meanwhile, for top-gradient neurons, when setting the activation shift range to ±2, the correlations in all models are above 0.95, which we consider the threshold for indicating the linear relationship. Our subsequent analysis all uses the top-gradient neurons within a shift range of ±2 by default.
@@
Neuron linearity. We then present a quantitative analysis of the prevalence of neuron linearity and the generality of these neurons across different prompts and Transformer layers. Specifically, we report the ratio of neurons exhibiting ‘linearity,’ defined as having correlations equal to or greater than 0.95 within a shift range of $\pm$2.666As there is no strict definition of linearity, we use the 0.95 as it indicates a strong linear relationship. To enhance the coverage of analysis results, we use 1000 prompts paired with 100 top-gradient neurons, conducting 100K neuron intervention experiments per PLM. [^7] We only chose 200 prompts and 100 neurons for Llama2-70B due to the large model size. The ratios of ‘linear’ neurons in BERTbase/large/wwm models are 0.963/0.955/0.980, respectively, and the ratios for Llama-7B/13B/70B are 0.914/0.917/0.977, indicating that a large portion of neurons exhibit linearity. Our analysis in § A.2 reveals that linear neurons are common across layers and prompts.
Neuron polarity. In the following discussions, we consider the direction of change in output probabilities in our numerical analysis. We denote neurons are positive/negative if increasing/decreasing their activations enhances the target output probabilities.
3 Neuron Empirical Gradient
@@ 3.1 NeurGrad
We quantify how important a neuron is in influencing the target token’s probability by the gradient of the linear relationship we term neuron empirical gradient. Empirical gradient provides a quantified measurement of neurons’ controllability: the magnitude indicates the intensity of control, while the polarity decides the controlling direction over model outputs. To calculate the neuron empirical gradient, we fit a zero-intercept linear regression between activation shifts and output probability changes acquired through neuron intervention. The regression coefficient is identified as the neuron empirical gradient, which requires extensive inferences for a specific neuron, prompt, and token.
To address this issue, we propose NeurGrad, inspired by the observation that computational gradients approximate empirical gradient magnitudes but fail to accurately capture neuron polarity, which is negatively correlated with activation signs.
$\bar{G_{E}}$ $=G_{C}\times-\operatorname{sign}(A),$ (1)
where $\bar{G_{E}}$, $A$, $G_{C}$, and $\operatorname{sign}(A)$ represents the estimated empirical gradient, activation, computational gradient, and sign of $A$ [^8] For neurons with zero activation, we assign an empirical gradient of zero as such cases are rare. On average, there are at most around 1000 zero-activation neurons in the PLMs. (1 for $A>0$ and -1 for $A<0$), respectively.
$G_{C}$ IG. NeurGrad
BERTlarge -.9307 .7360 .9998
BERTbase -.8909 .7167 .9958
BERTwwm -.8914 .8584 .9989
Llama2-7B .0115 .6728 .9769
Llama2-13B -.0113 .6964 .9641
Llama2-70B -.0391 n/a .7811
Table 1: Pearson correlations between the estimated and ground-truth empirical gradients using sampled neurons; memory cost precludes results of IG. for Llama2-70B.
Figure 3: Cumulative distribution of empirical gradient magnitudes as a function of the percentiles of absolute empirical gradients, sorted by descending order.
To validate NeurGrad’s effectiveness, we obtain ground-truth empirical gradients via neuron-wise intervention experiments. Then, we measure the Pearson correlation between ground-truth empirical gradients and NeurGrad-estimated gradients. Specifically, we collect empirical gradients of 1000 prompts, with 100 random neurons per prompt. The activation shift range is set to [-2, 2] according to § 2.2. We estimate the empirical gradients using three different methods: computational gradient ($G_{c}$), integrated gradients (IG.) used for identifying knowledge neurons (Dai et al., 2022) that intervene neuron in small step sizes multiple times to simulate the gradient, and NeurGrad.
Table 1 reports the Pearson correlations between estimated gradients and empirical gradients. The results indicate NeurGrad’s superiority in accurately measuring empirical gradients. NeurGrad is also much more efficient than IG. Regarding efficiency, calculating IG requires multiple iterations, each involving changes to neuron activations. In contrast, NeurGrad completes the calculation with just one inference pass, resulting in a computational cost nearly identical to that of computational gradients.
3.2 Understanding Neuron Controllability
The efficiency and precision of NeurGrad allow us to estimate empirical gradients of all neurons in LLMs. In this section, we leverage NeurGrad-estimated gradients to the following questions.
1). How are neurons distributed across layers? Figure 4 illustrates the means and variances of empirical gradient magnitudes across layers. The mean gradient magnitude reflects the intensity with which PLMs adjust output probabilities through neurons in a given layer, while the variance indicates how concentrated the effective neurons are within that layer. A positive correlation is observed between variances and means, [^9] The Pearson correlation between means and variances of neuron magnitudes across different layers are 0.88, 0.79, 0.87 for BERTbase/large/wwm, and 0.57, 0.51, 0.42 for Llama2-7B/13B/70B. suggesting that as PLMs increase the intensity of gradient activity in specific layers, they also focus more on a limited subset of neurons. In the shallow layers, gradients are generally low and dispersed across neurons, indicating that these layers perform broad linguistic processing, with many neurons contributing modestly to the final predictions. In contrast, deeper layers show increased gradient magnitudes and higher concentrations, reflecting more targeted and confident adjustments by the model. Specifically, in BERT models, a strong correlation is evident between layer depth and neuron controllability intensity, with deeper layers exhibiting larger gradient magnitudes. In contrast, Llama2 displays a distinct pattern: gradient magnitudes peak in the middle layers, decrease towards the deeper layers, and then unexpectedly increase at the final layers.

### v2

# Neuron Empirical Gradient: Discovering and Quantifying Neurons’ Global Linear Controllability

@@ Abstract
Although feed-forward neurons in pre-trained language models (PLMs) can store knowledge and their importance in influencing model outputs has been studied, existing work focuses on finding a limited set of neurons and analyzing their relative importance. However, the global quantitative role of activation values in shaping outputs remains unclear, hindering further advancements in applications like knowledge editing. Our study first investigates the numerical relationship between neuron activations and model output and discovers the global linear relationship between them through neuron interventions on a knowledge probing dataset. We refer to the gradient of this linear relationship as neuron empirical gradient (NEG), and introduce NeurGrad, an accurate and efficient method for computing NEG. NeurGrad enables quantitative analysis of all neurons in PLMs, advancing our understanding of neurons’ controllability. Furthermore, we explore NEG’s ability to represent language skills across diverse prompts via skill neuron probing. Experiments on MCEval8k, a multi-choice knowledge benchmark spanning various genres, validate NEG’s representational ability. The data and code are released. [^1] https://anonymous.4open.science/r/NeurGrad
@@
Discovering and Quantifying Neurons’ Global Linear Controllability
Xin Zhao    Zehui Jiang The University of Tokyo {xzhao,zjiang}@tkl.iis.u-tokyo.ac.jp                        Naoki Yoshinaga Institute of Industrial Science, The University of Tokyo ynaga@iis.u-tokyo.ac.jp
@@ 1 Introduction
Transformer (Vaswani et al., 2017)-based pretrained language models (PLMs) exhibit a remarkable ability to possess human knowledge, drawing significant attention to understand their internal mechanisms. While prior studies have highlighted the crucial role of feed-forward (FF) layers’ neurons in representing diverse knowledge, including factual knowledge (Dai et al., 2022; Yu and Ananiadou, 2024) and general language skills (Wang et al., 2022; Tan et al., 2024), they face several challenges. First, the existing methods primarily focus on ranking neurons to identify important ones (Dai et al., 2022; Meng et al., 2022; Yu and Ananiadou, 2024), yet they lack a direct and global quantitative measurement between neuron activations and model output, limiting their applicability to neuron-level adjustment applications, like knowledge editing Zhang et al. (2024). Second, the existing neuron importance-ranking methods are computationally expensive, requiring either multiple activation modification (Dai et al., 2022; Meng et al., 2022; Goldowsky-Dill et al., 2023) or extensive tensor calculation (Yu and Ananiadou, 2024), making inefficient to analyze all neurons on large models across diverse prompts, thereby hindering a global understanding of their roles.
@@
Figure 1: Overview of our contributions: i) observation of the neuron linearity, ii) an efficient method, NeurGrad, to quantify the linearity, iii) skill neuron probing on MCEval8K to confirm NEG captures language skills.
Initially, our study conducts a quantitative analysis to understand how neuron activations determine model outputs (RQ1). We investigate this by gradually modifying activations of randomly sampled neurons and observing the resulting changes in the probabilities of target tokens for correct knowledge (hereafter, output shift), using MyriadLAMA (Zhao et al., 2024), a factual knowledge probing dataset on nine PLMs, including large language models (LLMs) like Llama2-70B (§ 3). Notably, we discover that, within a certain range of activations, shifts in neuron activations (hereafter, activation shifts) have a linear relationship with the output shift. We call and quantify the gradient of this linear relationship as neuron empirical gradient (NEG), enabling quantitative neuron analysis.
Next, we ask: can we precisely adjust PLMs’ output probabilities by shifting neuron activations? (RQ2). As estimating NEG requires extensive inference over PLMs, we first introduce NeurGrad (§ 4), an accurate yet efficient method to calculate NEG for the single neuron, grounded in the empirical observation that computational gradients( 4.1) strongly correlate with NEG magnitudes but weakly with their directions. We validate its performance using MyriadLAMA by measuring their relative relationship to ground-truth NEGs. NeurGrad even shows superiority over baselines, including existing neuron-ranking methods (Dai et al., 2022; Yu and Ananiadou, 2024). Leveraging NeurGrad, we further investigate neuron controllability over multiple neurons through multi-neuron intervention (§ 5). We observe that NEG can be accumulated with multiple neurons, while the correlation diminishes as more neurons are involved or larger activation shifts are applied.
Finally, we investigate whether NEG can represent general language skills associated with diverse prompts rather than specific factual prompts (RQ3). To answer this, we conduct skill neuron probing that identifies neurons associated with language skills using NEGs as inputs. While prior studies conducted skill neuron probing, they focus on using neuron activation to build probers, leaving the representational ability of NEG unexamined (Wang et al., 2022; Song et al., 2024). Furthermore, as they only used multi-choice datasets with limited language skills, we introduce MCEval8K, a multi-choice knowledge evaluation benchmark spanning six genres and 22 tasks for comprehensive LLMs evaluation. Our experiment discovers NEG’s ability to represent diverse language skills, providing a basis for future language skill-oriented neuron-based model adjustments.
@@
  • We confirm that activation shifts are linearly correlated to output shifts within a specific activation shift range, referred to and quantified as neuron empirical gradients. (§ 3)
@@
  • We present NeurGrad, an efficient method for estimating neuron empirical gradients(§ 4), and conduct analysis to deepen the understanding of neuron controllability (§ 5).
  • We find NEG’s ability to express language skills by skill neuron probing (§ 6,§ 7).
@@
  • We built MCEval8K, a multi-choice benchmark spanning six knowledge genres and 22 language-understanding tasks (§ 6.2, § F).
2 Related Work
Neuron-level knowledge attribution methods. Existing studies built the connections between knowledge and neurons by measuring the importance scores of neurons to the model prediction on the target knowledge; some of them rely on causal observations of how knowledge changes with neuron modifications (Meng et al., 2022; Geva et al., 2021; Wang et al., 2024; Chen et al., 2023), while others use extensive tensor calculations to estimate neuron contributions (Geva et al., 2022; Lee et al., 2024; Yu and Ananiadou, 2024). However, these methods are computationally expensive, limiting their scalability for large-scale probing across diverse prompts in LLMs. Moreover, they often capture relative relationships (e.g., neuron rankings) rather than the direct, quantitative relationships between specific neurons and model outputs, restricting neurons’ utility in scenarios requiring precise, such as knowledge editing Zhang et al. (2024) and bias mitigation Gallegos et al. (2024) on LLMs. Although gradient-based approaches (Lundstrom et al., 2022; Dai et al., 2022) seek to integrate gradients through neuron intervention, they also suffer from high computational costs.
Skill neuron probing. Neurons in FF layers show the ability to convey specific skills so that using the neuron activations solely can tackle the language tasks, which these neurons are referred to as skill neurons (Wang et al., 2022; Song et al., 2024). Existing studies found neurons can express semantic skills like sentiment classification (Wang et al., 2022; Song et al., 2024) or complex skills, such as style transfer (Lai et al., 2024) and translation (Tan et al., 2024). However, previous research viewed neuron activations as knowledge indicators, and the representational ability of neuron gradients to language skills is not examined, hindering the application of neuron-level model adjustment.
3 Neuron Linearity to Model Output
This section empirically answers how neurons in PLMs’ FF layers influence model output by observing the resulting change in output tokens’ probabilities for fine-grained neuron-level interventions.
3.1 Neuron-level Intervention Experiment
@@
Models. To make the analysis result general, we experiment with masked and causal LMs of varied sizes and learning strategies. For masked LMs, we use three BERT (Vaswani et al., 2017; Devlin et al., 2019) models: BERTbase, BERTlarge, and BERTwwm. We construct masked prompts and let the model predict the masked token. For causal LMs, we examine diverse open-source LLMs, including instruction-tuned and pre-trained LLMs. The instruction-tuned LLMs are selected from two families: Llama2 (Touvron et al., 2023) with sizes of 7B, 13B, and 70B, and Phi3-mini (Abdin et al., 2024), a model with 3.8B parameters. The pre-trained LLMs are Llama2 models with 7B and 13B parameters. Following the zero-shot prompt setting in Zhao et al. (2024), we instruct them to generate single-token answers. See § E for model details.
@@
Dataset. We utilize a multi-prompt knowledge probing dataset, MyriadLAMA (Zhao et al., 2024), for neuron intervention. MyriadLAMA offers diverse prompts per fact, reducing the influence of specific linguistic expressions on probing results. We focus on single-token probing, where the target answer is represented by a single token. For each PLM, we randomly sample 1000 prompts from MyriadLAMA, where the model correctly predicts the target token. Due to differences in tokenizers, the probing prompts may vary across PLMs, and the results are not comparable across the PLMs.
@@
Neuron-wise intervention. We conduct a neuron-wise intervention to analyze how activation shift affects model outputs. Specifically, we alter the neuron activations within a range of [-10, 10] with a step size of 0.2 to observe the resulting changes in target token output probabilities. To establish a global observation over all neurons while minimizing computational cost, since assessing a single neuron’s effect on one token for one prompt requires 100 inferences, we conduct neuron interventions on randomly sampled neurons.
@@
Figure 2: Average absolute Corr between activation shifts and output shifts.
Result and Analysis To understand how output shifts respond to neuron activation shifts, we calculate the Pearson correlation (hereafter, Corr) between the activation shift ranges and the output shifts of the correct tokens, considering only the absolute Corr values. The absolute Corr is averaged over 10 prompts randomly sampled from 1000 prompts to reduce the computational cost, each with 1000 randomly sampled neurons given a specific shift range (x-axis). Figure 2 shows that all activation shifts and output shifts have a nearly linear, strong Corr, even with a broad range222$\pm$10 is an appropriate range considering the distribution of activations; see § B.1 for details. of $\pm$10. Figure 2 also indicates that activation shifts tend to show stronger Corr with output shifts at smaller shift ranges, consistent across all models. It indicates that the shifting neuron activation with a specific value can have a predictable result in output shifts.
Noted that as all PLMs show similar behaviors, we consider the BERT PLMs and three instruction-tuned Llama2 LLMs for the following analysis.
3.2 Neuron Linearity
Based on the above discussion, we ask: do neurons generally exhibit linearity to model output? We first define neurons as possessing linearity if their correlation (Corr) is at least 0.95333Since linearity lacks a strict definition, we use Corr¿=0.95 to indicate a strong linear relationship. within a shift range of $\pm$2 from the observation in Figure 2.
@@ 3.2 Neuron Linearity
We present a quantitative analysis of the prevalence of neuron linearity across different prompts and Transformer layers. Specifically, we report the ratio of neurons exhibiting linearity from 1000 prompts paired with 100 neurons. [^4] We only chose 200 prompts and 100 neurons for Llama2-70B due to the large model size. The ratios of ‘linear’ neurons in BERTbase/large/wwm models are 0.9565/0.8756/0.9564, respectively, and the ratios for Llama-7B/13B/70B are 0.9387/0.9677/0.9208, indicating that majority of neurons exhibit linearity. We analyze the generality of linear neurons across layers and prompts, revealing their widespread presence (§ A). Furthermore, we term the linearity direction as polarity, where neurons are positive/negative if increasing/decreasing their activations enhances the target output probabilities.
Neuron empirical gradient. We quantify the neurons’ linearity and polarity by the gradient of the linear relationship between activation shift and output shifts, which we term neuron empirical gradient (NEG). We calculate NEG as follows: we fit a zero-intercept linear regression between activation shifts and output shifts acquired through neuron intervention, and the regression coefficient is identified as the NEG for a specific neuron, prompt, and token.
Corr MAE
CG IG LPI NeurGrad CG IG LPI NeurGrad
BERTbase -.8909 .7360 - [^5] We follow code released in Yu and Ananiadou (2024) and only Llama2 LLMs are supported. .9998 6.1e-03 3.0e-03 - 2.6e-05
BERTlarge -.9307 .7167 - .9958 4.6e-03 2.1e-03 - 1.9e-04
BERTwwm -.8914 .8584 - .9989 4.5e-03 2.2e-03 - 2.3e-05
Llama2-7B .3020 .5383 .6469 .8136 1.7e-06 1.2e-06 1.9e-03 1.3e-06
Llama2-13B -.1973 .7261 .1141 .9965 2.1e-06 1.5e-06 4.3e-04 6.0e-08
Llama2-70B .0283 n/a [^6] Due to the high memory cost of IG and LPI, we preclude Llama2-70B experiments on these methods. n/a 1.000 5.9e-07 n/a n/a 2.1e-09
Avg. Runtime (Llama2-7B) 0.149s 19.349s 6.086s 0.161s Same as left
Table 1: Evaluation of NeurGrad and baselines in calculating NEGs, including two metrics: Corr and MAE.
4 NeurGrad for NEG Estimation
Efficiently and accurately computing NEG is crucial for quantitative analysis of neuron-level interpretability in PLMs. However, quantifying NEG through neuron-wise intervention is impractical due to the high computational cost. While prior studies proposed knowledge attribution methods that measure neurons’ importance in influencing model output, these methods require either extensive calculation or can only estimate the neurons’ relative importance but cannot directly estimate the NEG (Dai et al., 2022; Geva et al., 2022; Meng et al., 2022; Yu and Ananiadou, 2024).
@@ 4.1 NeurGrad
[…]
$\bar{G_{E}}$ $=\text{CG}\times\operatorname{sign}(A),$ (1)
where $\bar{G_{E}}$, $A$, and $\operatorname{sign}(A)$ represents the estimated NEG, activation, and sign of $A$ (1 for $A>0$ and -1 for $A<0$), respectively.
4.2 NEG Estimation Evaluation
We evaluate NeurGrad’s effectiveness in estimating NEG. We use the same setting as CG evaluation above to collect the ground-truth NEGs. Our experiment involves three baselines, including two gradient estimation methods, CG and integrated gradients (IG) (Dai et al., 2022), and one logit-based knowledge attribution method (Yu and Ananiadou, 2024). IG intervenes with neurons in small step sizes multiple times to simulate the NEG. The Log-Probability-Increase (LPI) (Yu and Ananiadou, 2024) identifies important neurons by estimating increased probabilities with specific neurons. [^8] We did not consider other causal-tracing-based methods (Meng et al., 2022) due to their high computational cost, which conflicts with our efficiency goal.
We evaluate how well these methods calculate NEG with two metrics: Corr and mean absolute error (MAE). Table 1 (left) reports the Corr between the estimated and ground-truth NEGs that can assess their ability to capture the neurons’ relative relationship, such as ordering. NeurGrad consistently achieves high Corr over the six PLMs, outperforming all other methods. Moreover, Table 1 (right) reports the Corr the MAE that quantifies the accuracy of NEGs’ calculation. It is observed that NeurGrad can largely reduce the estimation error to the true NEG, revealing its ability to precisely calculate NEGs. Finally, the average running time (Table 1 (bottom)) demonstrates NeurGrad’s efficiency compared to other methods. The time estimation uses Llama2-7B with an NVIDIA RTX A6000 GPU.
Figure 3: Comparison of neuron attribution methods in token probability enhancement. X-axis: activation shifts of selected neurons; Y-axis: average output shifts over 1000 factual prompts.
4.3 Knowledge Attribution Evaluation
Finally, we evaluate NeurGrad’s ability to locate important neurons. Specifically, for 1000 prompts, we identify the top-$K$ neurons ($K=1,4,16$) with the highest attribution scores using CG, IG, LPI, and NeurGrad. We enhance neuron activations by shifting positive neurons from 0.1 to 1 and negative neurons from -0.1 to -1, both in increments of 0.1. Figure 3 reports the output shifts of target tokens on Llama2-7b, with different activation shifts and top-k neurons selected from different attribution methods. Figure 3 demonstrates that NeurGrad outperforms baselines in attribution accuracy. This superiority stems from NeurGrad’s precise measurement of NEG and its inclusion of negative neurons in intervention, unlike IG and LPI, which only consider positive neurons despite their equal ratio to negative neurons (Table 2). See details in § C.
5 Understanding Neurons’ Controllability
This section deepens our understanding of neuron controllability: the ability to precisely adjust PLM output probabilities by modifying neuron activations, using NEGs estimated by NeurGrad.